# Computation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v2

@@ 4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
As to the first one, suppose that there is a $K\in\mathcal{Q}(L)$ which belongs to $\psi\circ\phi(\mathcal{A})$ but not to $\mathcal{A}$. Then $K\in\mathcal{Q}(L)\setminus\mathcal{A}$, which is Scott open in $\mathcal{Q}(L)$ and further an open subset of $\mathcal{Q}(L)$ with respect to the upper Vietoris topology by $(\mathrm{i})$. Thus there exists a $U\in\sigma(L)$ such that $K\in\Box U\subseteq\mathcal{Q}(L)\setminus\mathcal{A}$, which implies $K\cap(L\setminus U)=\emptyset$. This will contradict the fact that $K\in\psi\circ\phi(\mathcal{A})$ if we could show $L\setminus U\in\phi(\mathcal{A})$. Assume $(L\setminus U)\cap K_{0}=\emptyset$ for some $K_{0}\in\mathcal{A}$. Then $K_{0}\in\Box U$, thus $K_{0}\notin\mathcal{A}$, a contradiction. So $L\setminus U\in\phi(\mathcal{A})$. Therefore, $\psi\circ\phi(\mathcal{A})\subseteq\mathcal{A}$.
Regarding the second, for the sake of a contradiction, we assume that there is an $A\in\phi\circ\psi(\mathcal{K})$ but $A\notin\mathcal{K}$. It suggests that each $C\in\mathcal{K}$ is not contained in $A$, that is to say, $(L\setminus A)\cap C\neq\emptyset$ for all $C\in\mathcal{K}$. Since $L$ has the property $(\mathrm{KC})$, we can find a $K_{0}\subseteq L\setminus A$ intersecting each $C$ in $\mathcal{K}$. Then $K_{0}\in\psi(\mathcal{K})$ and $K_{0}\cap A=\emptyset$, which contradicts that $A\in\phi\circ\psi(\mathcal{K})$. Thus $A\in\mathcal{K}$, which obviously leads to a contradiction. So our assumption is wrong and $\phi\circ\psi(\mathcal{K})\subseteq\mathcal{K}$ holds. ∎
It was Schalk who proved that for a locally compact sober space $X$, the upper Vietoris topology and the Scott topology on $Q(X)$ coincide (see [20, Lemma 7.26]), meanwhile, [3, Proposition 5.4] reveals that each locally compact space is consonant. Therefore, combining Lemma 4.5 and Theorem 4.8, we further draw the conclusion in the following.
@@ 5 The double power well-filtered dcpo
$\mathcal{T}$$\mathcal{TS}$$\mathcal{T}\eta^{\mathcal{S}}$$\mathcal{ST}$$\eta^{\mathcal{S}}_{\mathcal{T}}$$l$$\mathcal{S}$$\mathcal{TS}$$\mathcal{ST}$$\mathcal{S}\eta^{\mathcal{T}}$$\eta^{\mathcal{T}}_{\mathcal{S}}$$l$$\mathcal{TSS}$$\mathcal{STS}$$\mathcal{SST}$$l_{\mathcal{S}}$$\mathcal{S}l$$\mathcal{TS}$$\mathcal{ST}$$T\mu^{\mathcal{S}}$$\mu^{\mathcal{S}}_{\mathcal{T}}$$l$$\mathcal{TTS}$$\mathcal{TST}$$\mathcal{STT}$$\mathcal{T}l$$l_{\mathcal{T}}$$\mathcal{TS}$$\mathcal{ST}$$\mu^{\mathcal{T}}_{\mathcal{S}}$$\mathcal{S}\mu^{\mathcal{T}}$$l$
@@
$Id$$\mathcal{T}$$\mathcal{T}^{\prime}$$\eta$$\eta^{\prime}$$\lambda$$\mathcal{TT}$$\mathcal{T^{\prime}T}$$\mathcal{T^{\prime}T^{\prime}}$$\mathcal{T}$$\mathcal{T}^{\prime}$$\lambda_{\mathcal{T}}$$\mathcal{T}^{\prime}\lambda$$\mu$$\mu^{\prime}$$\lambda$
@@
      $(\mathcal{TS},\eta^{\mathcal{T}}\eta^{\mathcal{S}},\mu)$ is a monad;
@@
$\phi$ and $\psi$ given in Definition 4.1 are distributive laws of $\mathcal{Q}$ over $\mathcal{H}$ and of $\mathcal{H}$ over $\mathcal{Q}$, respectively.
@@
It is a routine to check by their definitions. ∎
Based on the above results, we conclude that both $\mathcal{H}\mathcal{Q}$ and $\mathcal{Q}\mathcal{H}$ are monads on $\mathbf{WF}$. Naturally, the problem that needs further consideration is whether the $\mathcal{Q}\mathcal{H}$-algebras (or $\mathcal{H}\mathcal{Q}$-algebras) are related to the $\mathcal{Q}$-algebras and the $\mathcal{H}$-algebras. In the following, we will take the monad $\mathcal{Q}\mathcal{H}$ as the representative for an investigation.
@@
$\gamma:$$Id$$\mathcal{Q}$$\mathcal{H}$$\mathcal{Q}\mathcal{H}$$\theta$$\eta$$\mathcal{Q}\eta$$\theta_{\mathcal{H}}$$\rho:$$\mathcal{Q}\mathcal{H}\mathcal{Q}\mathcal{H}$$\mathcal{Q}\mathcal{Q}\mathcal{H}\mathcal{H}$$\mathcal{Q}\mathcal{H}\mathcal{H}$$\mathcal{Q}\mathcal{Q}\mathcal{H}$$\mathcal{Q}\mathcal{H}$$\mathcal{Q}\phi_{\mathcal{H}}$$\iota_{\mathcal{H}\mathcal{H}}$$\mathcal{Q}\mathcal{Q}\mu$$\mathcal{Q}\mu$$\iota_{\mathcal{H}}$
@@
([8]) Let $\mathbf{C}$ be a poset-enriched category. A monad $\mathcal{T}=(\mathcal{T},\eta,\mu)$ on $\mathbf{C}$ is called a right $\rm{KZ}$-monad if $\mathcal{T}$ is a poset-functor and subjects to $\eta_{TX}\leq T\eta_{X}$ for all $X\in\mathbf{C}$.
Left $\rm{KZ}$-monads are defined poset-dually, by reversing the inequality.
@@
Theorem 5.7.
([8]) Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a right $\rm{KZ}$-monad in a poset-enriched category $\mathbf{C}$. Then the following statements are equivalent for any object $C\in\mathbf{C}$.
  1. 1.
    $C$ is a $T$-algebra.
  2. 2.
    $C$ is injective over the right $T$-embeddings.
Note that a right $T$-embedding mentioned in the above theorem is an arrow in $\mathbf{C}$ satisfying that $Tf$ has a left adjoint $h$ (i.e., $h\circ Tf\leq id$ and $Tf\circ h\geq id$), and $h\circ Tf=id$.
@@
The definitions of the two $\rm{KZ}$-monads make the verification trivial. ∎
Lemma 5.10.
@@
    $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra.
@@
    $\alpha$ is a $\mathcal{Q}$-algebra homomorphism and $\alpha\circ\phi$ is an $\mathcal{H}$-algebra homomorphism.

### v3

@@ 4 The commutativity between $\mathcal{H}$ and $\mathcal{Q}$
As to the first one, suppose that there is a $K\in\mathcal{Q}(L)$ belonging to $\psi\circ\phi(\mathcal{A})$ but not to $\mathcal{A}$. Then $K\in\mathcal{Q}(L)\setminus\mathcal{A}$, which is Scott open in $\mathcal{Q}(L)$ and further an open subset of $\mathcal{Q}(L)$ with respect to the upper Vietoris topology by $(\mathrm{i})$. Thus there exists a $U\in\sigma(L)$ such that $K\in\Box U\subseteq\mathcal{Q}(L)\setminus\mathcal{A}$, which implies $K\cap(L\setminus U)=\emptyset$. This will contradict the fact that $K\in\psi\circ\phi(\mathcal{A})$ if we could show $L\setminus U\in\phi(\mathcal{A})$. Assume $(L\setminus U)\cap K_{0}=\emptyset$ for some $K_{0}\in\mathcal{A}$. Then $K_{0}\in\Box U$, thus $K_{0}\notin\mathcal{A}$, a contradiction. So $L\setminus U\in\phi(\mathcal{A})$. Therefore, $\psi\circ\phi(\mathcal{A})\subseteq\mathcal{A}$.
Regarding the second, for the sake of a contradiction, we assume that there is an $A\in\phi\circ\psi(\mathcal{K})$ but $A\notin\mathcal{K}$. It suggests that each $C\in\mathcal{K}$ is not contained in $A$, that is to say, $(L\setminus A)\cap C\neq\emptyset$ for all $C\in\mathcal{K}$. Since $L$ has the property $(\mathrm{KC})$, we can find a $K_{0}\subseteq L\setminus A$ intersecting each $C$ in $\mathcal{K}$. Then $K_{0}\in\psi(\mathcal{K})$ and $K_{0}\cap A=\emptyset$, which contradicts that $A\in\phi\circ\psi(\mathcal{K})$. Thus $A\in\mathcal{K}$, which obviously leads to a contradiction. So our assumption is wrong and $\phi\circ\psi(\mathcal{K})\subseteq\mathcal{K}$ holds.
It was Schalk who proved that for a locally compact sober space $X$, the upper Vietoris topology and the Scott topology on $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{Q}(X)$} coincide (see [21, Lemma 7.26]), meanwhile, [3, Proposition 5.4] reveals that each locally compact space is consonant. Therefore, combining Lemma 4.6 and Theorem 4.8, we further draw the conclusion in the following.
@@ 5 The double power well-filtered dcpo
[…]
@@
[…]
@@
      $(\mathcal{TS},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\eta_{\mathcal{S}}^{\mathcal{T}}\eta^{\mathcal{S}}},\mu)$ is a monad;
@@
  3. 3.
    Liftings of $\tilde{\mathcal{S}}$ of the monad $\mathcal{S}$ into $\mathbf{C}^{\mathcal{T}}$.
From the proposition presented in Section 2 of Beck [2], a more general result concerning the morphisms of monads was established by Fritz et al.
Lemma 5.4.
@@
$\phi$ : $\mathcal{H}\mathcal{Q}\rightarrow\mathcal{Q}\mathcal{H}$ defined in Definition 4.1 is a distributive law of $\mathcal{Q}$ over $\mathcal{H}$.
@@
$\bf{Claim~1:}$ $\phi$ is a natural transformation.
Assume $L$ and $M$ are well-filtered dcpos and $f:L\rightarrow M$ is a Scott-continuous mapping. We need to show the following diagram commutes,
$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\phi_{L}$$\mathcal{HQ}(M)$$\mathcal{QH}(M)$$\mathcal{HQ}(f)$$\mathcal{QH}(f)$$\phi_{M}$
i.e., for any $\mathcal{A}\in\mathcal{HQ}(L)$, $\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})=\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$, where
$\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})={\uparrow}\{cl(f(A)):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}$,
$\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})=\{B\in\mathcal{H}(M):\forall N\in cl\{{\uparrow}f(K):K\in\mathcal{A}\},B\cap N\neq\emptyset\}$.
First, suppose that there exists a $cl(f(A))\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ but $cl(f(A))\notin\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$, that is, we can find a $N_{0}\in cl\{{\uparrow}f(K):K\in\mathcal{A}\}$ satisfying $cl(f(A))\cap N_{0}=\emptyset$. From which we can see that $N_{0}\in\Box(M\setminus cl(f(A)))$. Since the well-filteredness of $M$ implies that $\Box(M\setminus cl(f(A)))\in\sigma(\mathcal{Q}(M))$, $\Box(M\setminus cl(f(A)))\cap\{{\uparrow}f(K):K\in\mathcal{A}\}\neq\emptyset$. So there is a $K_{0}\in\mathcal{A}$ such that ${\uparrow}f(K_{0})\subseteq M\setminus cl(f(A))$, that is to say, ${\uparrow}f(K_{0})\cap cl(f(A))=\emptyset$. Then $K_{0}\cap A=\emptyset$, which obviously contradicts the fact that $A$ meets every element in $\mathcal{A}$. Thus, $\{cl(f(A)):\forall K\in\mathcal{A},A\cap K\neq\emptyset\}\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$; furthermore, $\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})\subseteq\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$.
Now we prove the reverse inclusion. Let $B\in\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})$. Then for each $K\in\mathcal{A}$, $B\cap{\uparrow}f(K)\neq\emptyset$, combing with the downness of $B$, which further indicates that $B\cap f(K)\neq\emptyset$. We pick $x_{K}\in K$ such that $f(x_{K})\in B\cap f(K)$ and set $A_{0}=cl(\{x_{K}:K\in\mathcal{A}\})$. Then $cl(f(A_{0}))=cl(f(\{x_{K}:K\in\mathcal{A}\}))=cl(\{f(x_{K}):K\in\mathcal{A}\})\subseteq B$ by the closeness of $B$. Meanwhile, $A_{0}\cap K\neq\emptyset$ for each $K\in\mathcal{A}$. So $B\in\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$, then, $\phi_{M}\circ\mathcal{HQ}(f)(\mathcal{A})\subseteq\mathcal{QH}(f)\circ\phi_{L}(\mathcal{A})$ holds.
$\bf{Claim~2:}$ The first diagram in Definition 5.1 commutes.
Concretely, given a well-filtered dcpo $L$, we need to demonstrate that the following diagram commutes.
$\mathcal{H}(L)$$\mathcal{HQ}(L)$$\mathcal{H}\theta_{L}$$\mathcal{QH}(L)$$\theta_{\mathcal{H}(L)}$$\phi_{L}$
Let $A\in\mathcal{H}(L)$. Then $\theta_{\mathcal{H}(L)}(A)={\uparrow}\{A\}$ and
$\phi_{L}\circ\mathcal{H}\theta_{L}(A)=\{B\in\mathcal{H}(L):\forall K\in cl(\{{\uparrow}x:x\in A\}):B\cap K\neq\emptyset\}$.
First, suppose that there is a $K\in cl(\{{\uparrow}x:x\in A\})$ such that $A\cap K=\emptyset$, then $K\subseteq L\setminus A$, i.e., $K\in\Box(L\setminus A)$. By the well-filteredness of $L$, we know $\Box(L\setminus A)\in\sigma(\mathcal{Q}(L))$. So $\Box(L\setminus A)\cap\{{\uparrow}x:x\in A\}\neq\emptyset$, that is to say, we can find an $x\in A$ such that ${\uparrow}x\subseteq L\setminus A$. Thus $x\notin A$, clearly, which is a contradiction. So for any $K\in cl(\{{\uparrow}x:x\in A\}),A\cap K\neq\emptyset$, which implies that $A\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$; hence, $\theta_{\mathcal{H}(L)}(A)={\uparrow}\{A\}\subseteq\phi_{L}\circ\mathcal{H}\theta_{L}(A)$.
Now consider the reverse inclusion. For any $B\in\phi_{L}\circ\mathcal{H}\theta_{L}(A)$, we have $B\cap{\uparrow}x\neq\emptyset,\forall x\in A$. It means that each $x\in A$ belongs to $B$. Thus $A\subseteq B$, equivalently, $B\in{\uparrow}\{A\}$. By the arbitrariness of $B$, we conclude that $\phi_{L}\circ\mathcal{H}\theta_{L}(A)\subseteq\theta_{\mathcal{H}(L)}(A)$.
$\bf{Claim~3:}$ The second diagram in Definition 5.1 commutes.
Specifically, given a well-filtered dcpo $L$, we need to prove that the following diagram commutes,
$\mathcal{Q}(L)$$\mathcal{HQ}(L)$$\eta_{\mathcal{Q}(L)}$$\mathcal{QH}(L)$$\mathcal{Q}\eta_{L}$$\phi_{L}$
i.e., to verify $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)=\mathcal{Q}\eta_{L}(K)$ for any $K\in\mathcal{Q}(L)$, where $\mathcal{Q}\eta_{L}(K)={\uparrow}\{{\downarrow}x:x\in K\}$ and
$\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)=\{A\in\mathcal{H}(L):\forall M\in{\downarrow}\{K\}:A\cap M\neq\emptyset\}$.
Let $A$ be an arbitrary element in $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$. Then $A\cap K\neq\emptyset$, and thus, there is an $x\in A\cap K$. By the closedness of $A$, we have ${\downarrow}x\subseteq A$, i.e., $A\in{\uparrow}\{{\downarrow}x\}$. So $A\in\mathcal{Q}\eta_{L}(K)$. The arbitrariness of $A$ guarantees that $\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)\subseteq\mathcal{Q}\eta_{L}(K)$.
Now focus on each $x\in K$, we have ${\downarrow}x\cap K\neq\emptyset$. So for any $M\in{\downarrow}\{K\}$, i.e., $K\subseteq M$, ${\downarrow}x\cap M$ always holds. This means $\{{\downarrow}x:x\in K\}\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$; thus, $\mathcal{Q}\eta_{L}(K)\subseteq\phi_{L}\circ\eta_{\mathcal{Q}(L)}(K)$.
$\bf{Claim~4:}$ The third diagram in Definition 5.1 commutes, i.e., the following one commutes:
$\mathcal{HQQ}(L)$$\mathcal{QHQ}(L)$$\mathcal{QQH}(L)$$\phi_{\mathcal{Q}(L)}$$\mathcal{Q}\phi_{L}$$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\mathcal{H}\tau_{L}$$\tau_{\mathcal{H}(L)}$$\phi_{L}$
For any $\mathbf{A}\in\mathcal{HQQ}(L)$, by computation, we have
$\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})=\bigcup{\uparrow}\{\phi_{L}(\mathcal{A}):\forall\mathcal{K}\in\mathbf{A},\mathcal{A}\cap\mathcal{K}\neq\emptyset\}$,
$\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})=\{A\in\mathcal{H}(L):\forall K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}),A\cap K\neq\emptyset\}$.
Consider $\phi_{L}(\mathcal{A})=\{B\in\mathcal{H}(L):\forall M\in\mathcal{A},B\cap M\neq\emptyset\}$, where $\mathcal{A}$ meets each element of $\mathbf{A}$. Suppose that there exists a $B\in\mathcal{H}(L)$ that belongs to $\phi_{L}(\mathcal{A})$ but not to $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$, i.e., there is a $K\in cl(\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\})$ satisfying that $B\cap K=\emptyset$. It further implies that $K\in\Box(L\setminus B)$. Applying the well-filteredness of $L$ again, we obtain that $\Box(L\setminus B)$ is a Scott open subset of $\mathcal{Q}(L)$; thus, $\Box(L\setminus B)\cap\{\bigcup\mathcal{K}:\mathcal{K}\in\mathbf{A}\}\neq\emptyset$. So we can find a $\mathcal{K}\in\mathbf{A}$ satisfying $\bigcup\mathcal{K}\subseteq L\setminus B$. From which we say that for any $C\in\mathcal{K}$, $C\cap B=\emptyset$. Combining with the fact that $B\in\phi_{L}(\mathcal{A})$, one can conclude that every $C\in\mathcal{K}$ is not in $\mathcal{A}$, that is to say, $\mathcal{K}\cap\mathcal{A}=\emptyset$. However, $\mathcal{A}$ meets every element of $\mathbf{A}$, a contradiction. Thus, each $\phi_{L}(\mathcal{A})$ is contained in $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$; accordingly, $\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$.
Take an arbitrary element $A\in\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})$, then for each $\mathcal{K}\in\mathbf{A},A\cap\bigcup\mathcal{K}\neq\emptyset$. So there is an $a_{\mathcal{K}}\in A\cap\bigcup\mathcal{K}$, which further indicates that ${\uparrow}a_{\mathcal{K}}\subseteq K$, i.e., $K\leq{\uparrow}a_{\mathcal{K}}$ for some $K\in\mathcal{K}$. Since $\mathcal{K}$ is an upper set, ${\uparrow}a_{\mathcal{K}}\in\mathcal{K},\forall\mathcal{K}\in\mathbf{A}$. We set $C=cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$. Then under the mapping $\mathcal{H}\theta_{L}$, we can obtain a Scott closed subset $cl(\{{\uparrow}c:c\in C\})$ of $\mathcal{Q}(L)$. Next, we verify that $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ and $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$. First, if there is an $M\in cl(\{{\uparrow}c:c\in C\})$ such that $A\cap M=\emptyset$, then $M\in\Box(L\setminus A)$. Since the well-filteredness of $L$ implies the Scott openness of $\Box(L\setminus A)$, $\Box(L\setminus A)\cap\{{\uparrow}c:c\in C\}\neq\emptyset$. That is, there exists a $c\in C$ such that ${\uparrow}c\subseteq L\setminus A$, then, $c\in L\setminus A$. As $L\setminus A\in\sigma(L)$ and $c\in cl\{a_{\mathcal{K}}:\mathcal{K}\in\mathbf{A}\}$, we can find an $a_{\mathcal{K}}$ such that $a_{\mathcal{K}}\in L\setminus A$. i.e., $a_{\mathcal{K}}\notin A$, obviously, this leads to a contradiction. Thus, $A\in\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$. Then, since for each $a_{\mathcal{K}}\in C$, ${\uparrow}a_{\mathcal{K}}\in cl(\{{\uparrow}c:c\in C\})$ and ${\uparrow}a_{\mathcal{K}}\in\mathcal{K}$, $cl(\{{\uparrow}c:c\in C\})\cap\mathcal{K}\neq\emptyset,\forall\mathcal{K}\in\mathbf{A}$. So $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))\in\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$. As a conclusion, we find an element $\phi_{L}(cl(\{{\uparrow}c:c\in C\}))$ in $\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$ such that $A$ belongs to it, by the arbitrariness of $A$, we get that $\phi_{L}\circ\mathcal{H}\tau_{L}(\mathbf{A})\subseteq\tau_{\mathcal{H}(L)}\circ\mathcal{Q}\phi_{L}\circ\phi_{\mathcal{Q}(L)}(\mathbf{A})$ holds.
$\bf{Claim~5:}$ The forth diagram in Definition 5.1 commutes, i.e., the following one commutes:
$\mathcal{HHQ}(L)$$\mathcal{HQH}(L)$$\mathcal{QHH}(L)$$\mathcal{H}\phi_{L}$$\phi_{\mathcal{H}(L)}$$\mathcal{HQ}(L)$$\mathcal{QH}(L)$$\mu_{\mathcal{Q}(L)}$$\mathcal{Q}\mu_{L}$$\phi_{L}$
Let $\mathbf{A}$ be an arbitrary element in $\mathcal{HHQ}(L)$. By computation, we have
$\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})=\{A\in\mathcal{H}(L):\forall K\in\bigcup\mathbf{A},A\cap K\neq\emptyset\}$,
$\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})={\uparrow}\{\bigcup\mathcal{C}:\mathcal{C}\in\mathcal{HH}(L),\forall\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\cap\mathcal{K}\neq\emptyset\}$.
For each $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$, we have ${\downarrow}\{A\}\in\mathcal{HH}(L)$ and $A=\bigcup{\downarrow}\{A\}$. Suppose that there exists a $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\})$ such that ${\downarrow}\{A\}\cap\mathcal{K}=\emptyset$, then $\mathcal{K}\in\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$. By the well-filteredness of $\mathcal{H}(L)$, we know $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})$ is a Scoot open subset of $\mathcal{QH}(L)$. Therefore, $\Box(\mathcal{H}(L)\setminus{\downarrow}\{A\})\cap\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}\neq\emptyset$. So there is an $\mathcal{A}\in\mathbf{A}$ such that $\phi_{L}(\mathcal{A})\subseteq\mathcal{H}(L)\setminus{\downarrow}\{A\}$. This further indicates that $\phi_{L}(\mathcal{A})\cap{\downarrow}\{A\}=\emptyset$, i.e., $A\notin\phi_{L}(\mathcal{A})$. From which we derive that $A\cap K=\emptyset$ for some $K\in\mathcal{A}$. Note that $\mathcal{A}\in\mathbf{A}$ implies $K\in\bigcup\mathbf{A}$. As a result, clearly, $A\cap K=\emptyset$ contradicts the fact that $A\in\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$. So $A=\bigcup{\downarrow}\{A\}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$, and hence, $\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})\subseteq\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$.
Pick an arbitrary element $\bigcup\mathcal{C}$, where $\mathcal{C}\in\mathcal{HH}(L)$ and for each $\mathcal{K}\in cl(\{\phi_{L}(\mathcal{A}):\mathcal{A}\in\mathbf{A}\}),\mathcal{C}\cap\mathcal{K}\neq\emptyset$. Assume that there exists a $K\in\bigcup\mathbf{A}$, i.e., $K\in\mathcal{A}$ for some $\mathcal{A}\in\mathbf{A}$, such that $\bigcup\mathcal{C}\cap K=\emptyset$. Then for each $C\in\mathcal{C}$, $C\cap K=\emptyset$, which immediately means that $C\notin\phi_{L}(\mathcal{A}),\forall C\in\mathcal{C}$, in other words, $\mathcal{C}\cap\phi_{L}(\mathcal{A})=\emptyset$. This contradicts $\bigcup\mathcal{C}\in\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})$. So $\mathcal{Q}\mu_{L}\circ\phi_{\mathcal{H}(L)}\circ\mathcal{H}\phi_{L}(\mathbf{A})\subseteq\phi_{L}\circ\mu_{\mathcal{Q}(L)}(\mathbf{A})$.
To sum up, $\phi$ defined in Definition 4.1 is a distributive law of $\mathcal{Q}$ over $\mathcal{H}$.
From the above results, we derive that $\mathcal{Q}\mathcal{H}$ is a monad on $\mathbf{WF}$. Notably, Steven Vickers studied the Eilenberg-Moore algebras of the double powerlocale $P_{U}P_{L}$ (where $P_{U}$ and $P_{L}$ denote the upper and lower powerlocales, which commute up to isomorphisms via a distributive law) and it is concluded that $P_{U}P_{L}$-algebras correspond to localic frames [24]. Motivated by this work, we now investigate the algebraic structures of $\mathcal{Q}\mathcal{H}$-algebras.
@@
[…]
@@
([8]) Let $\mathbf{C}$ be a poset-enriched category. A monad $\mathcal{T}=(\mathcal{T},\eta,\mu)$ on $\mathbf{C}$ is called a right $\rm{KZ}$-monad if $\mathcal{T}$ is a poset-functor and satisfies $\eta_{TX}\leq T\eta_{X}$ for all $X\in\mathbf{C}$. Left $\rm{KZ}$-monads are defined poset-dually, by reversing the inequality.
@@
Proposition 5.8.
[8] Let $\mathcal{T}=(\mathcal{T},\eta,\mu)$ be a $\rm{KZ}$-monad on the category $\mathbf{C}$. Then by $(\mathrm{KZ}_{1})$, every object has at most one structure map.
@@
The definitions of the two $\rm{KZ}$-monads make the proof trivial.
@@
    $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ and $(L,\alpha^{\mathcal{Q}})$ with $\alpha^{\mathcal{Q}}=\alpha\circ\mathcal{Q}\eta_{L}$ are $\mathcal{Q}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism.
@@
    $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})$ and $(L,\alpha^{\mathcal{H}})$ with $\alpha^{\mathcal{H}}=\alpha\circ\theta_{\mathcal{H}(L)}$ are $\mathcal{H}$-algebras and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{H}})$ is an $\mathcal{H}$-algebra homomorphism.